OpenStudy (anonymous):

The variable z varies jointly as the square root of x and the cube root of y. If z=108.0 when x=100 and y=64, find z when x=9 and y=27.

OpenStudy (anonymous):

How do I set this up and do it?

OpenStudy (anonymous):

@ganeshie8 @jim_thompson5910 @phi @thomaster

OpenStudy (anonymous):

@hartnn @Luigi0210 @Preetha @bibby

OpenStudy (anonymous):

@Destinymasha @kropot72 @ranga @beccaboo333 @sourwing @someonepls

OpenStudy (anonymous):

So, I've personally never heard it said this way, but "Z varies jointly with..." implies: \[Z = C \sqrt{x}\sqrt[3]{y}\] where C is some constant. We are told that when x = 100 and y = 64, that z = 108. We plug these values in and solve for C. We can then find Z for the given x and y.

OpenStudy (anonymous):

So \[108=C \sqrt{100}\sqrt[3]{64}\]

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

So that's 100=C(40) Which is C=2.5, amirite?

OpenStudy (anonymous):

108 = 40 C C = 2.7

OpenStudy (anonymous):

Oh my bad, I wrote it incorrectly

OpenStudy (anonymous):

So \[z=2.7\sqrt{9}\sqrt[3]{27}\]

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

z=24.3?

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

Thanks! Can you help me with another one?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Give me a bit to write it down

OpenStudy (anonymous):

The kinetic energy K of a moving object varies jointly with its mass m and the square of the velocity v. If an object weighing 17 kilograms and moving with a velocity of 10 meters per second has a kinetic energy of 850 Joules, find its kinetic energy when the velocity is 22 meters per second.

OpenStudy (anonymous):

There we go. Let me try to set it up real quick to see if I learned anything :D

OpenStudy (anonymous):

\[K=Cm \sqrt{v}\]

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

Close. It varies with the SQUARE of the velocity. Not the square root.

OpenStudy (anonymous):

Ohh \[K=Cmv ^{2}\]

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

:) So \[850=C(17)(10^{2})\]

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

850=C(170) 850/170=C C=5

OpenStudy (anonymous):

10 * 10 = 100 100* 17 = 1700

OpenStudy (anonymous):

My god, it's always the silly little mistakes that get me

OpenStudy (anonymous):

Haha, I know the feeling :)

OpenStudy (anonymous):

850/1700=0.5

OpenStudy (anonymous):

Soooooo \[K=0.5(17)(22^{2})\]

OpenStudy (anonymous):

K=0.5(8228)

OpenStudy (anonymous):

K=4114 is my final answer

OpenStudy (anonymous):

That's it :)

OpenStudy (anonymous):

Thank you so much! I got these last two questions right

OpenStudy (anonymous):

Outstanding! Any time :)